Find the solution to the system of equations given below by using linear combination:
8x + 4y = 48
4x – 2y = 40
A. (8, -4)
B.(11, 2)
C.(-8, 24)
D. (8, 4)

Answer by checkley79(3341)

8x + 4y = 48
4x – 2y = 40 MULTIPLY BY 2 & ADD
8X-4Y=80
8X+4Y=48
------------------
16X=128
X=128/16
X=8 ANS.
8*8+4Y=48
64+4Y=48
4Y=48-64
4Y=-16
Y=-4 ANS.
PROOF:
4*8-2*-4=40
32+8=40
40=40
ANSWER A(8,-4)